According to fiscal policy principles, a tax increase will OA. OB. OC. Increase spending and increase aggregate demand Reduce spending and reduce aggregate demand Reduce spending and increase aggregate demand D Increase spending and reduce aggregate demand. What is the weighted-average cost of capital for a firm using 65% common equity with a return of 15%, 25% debt with a return of 6%, 10% preferred stock with a return of 10%, and a tax rate of 35%? OA, OB. OC. OD. 11.725% 10.333% 11.275% 12.250% An issuer is preparing to file its annual report prior to adopting a code of ethics for its senior financial officers. What action, if any, must the issuer take to comply with the Sarbanes-Oxley Act of 2002? OA. Adopt a code of ethics for senior financial officers within 60 days of filing OB. Disclose in the annual report the lack of the code of ethics and the reason(s). OC. No action. OD. File the annual report after adopting a code of ethics for senior financial officers
Answer #1

Solution 1:

As per fiscal policy principle, A tax increase will result in reduce spending and reduce aggregate demand.It shifts the demand curve to left side.

Hence choice B seems to be correct.

Solution 2:

Cost of Equity (Ke) = 15%

Cost of debt = Interest rate (1-Tax) = 6% (1-0.35) = 3.9%

cost of preferred stock (Kp) = 10%

Weighted Average cost of capital = Ke*We + Kd*Wd + Kp*Wp

= (0.15*65%) + (0.039 * 25%) + (0.10*10%) = 11.725%